Can someone please, please help me with this problem:
What amount of cash would result at the end of one year, if $17,000 is invested today and the rate of return is 10%?
a. $17,000
b. $18,530
c. $18,700
d. none of the above

Answer by stanbon(75887)

What amount of cash would result at the end of one year, if $17,000 is invested today and the rate of return is 10%?
------------
A = P(1 + rt)
Amount = 17000(1 + 0.1*1)
Amount = 17000(1.1) = 18.700
================================
Cheers,
Stan H.